OpenStudy (anonymous):

verify 1-sin^2(2x)=(1-2sin^2(x))^2

OpenStudy (anonymous):

\[ 1-\sin^2(2x) \]Use the fact that \(\sin^2(y) + \cos^2(y) = 1\). If we let \(y=2x\) then we would say \(\sin^2(2x) + \cos^2(2x) = 1\).\[ [\sin^2(2x) + \cos^2(2x)]-\sin^2(2x) = \cos^2(2x) \]

OpenStudy (anonymous):

Then there is a double angle formula: \[ \cos(2x) = 1-2\sin^2(x) \]